OpenStudy (anonymous):

-13=3y-7 need help plzzzz

OpenStudy (anonymous):

2 or -2. my head is fuzzy

OpenStudy (anonymous):

\[-13=3y-7\]Move all terms containing y to the left, all other terms to the right. Subtract -3y from both sides. \[-13-3y = -7\] Add 13 to both sides. \[-3y = -7 +13\] Combine like terms. 13 -7 is 6. \[-3y =6\] Divide each side by -3. \[y = \frac{6}{-3}\] \[y = -2\]

OpenStudy (anonymous):

-13+7/3